OpenStudy (anonymous):

find the missing side of special right triangles. leave answer as radicals in simplest form.

OpenStudy (anonymous):



OpenStudy (anonymous):

use sine law..

OpenStudy (anonymous):

but i need it in radical form?!

OpenStudy (anonymous):

\[\sin(60)=\frac{7}{x}\] \[\frac{\sqrt{3}}{2}=\frac{7}{x}\] solve the ratio for \(x\)

OpenStudy (anonymous):

oh you need \(x\) and \(y\) ?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

i need x and y and in radical form! thanks so much ur so kind;)

OpenStudy (anonymous):

http://www.freemathhelp.com/triangle-30-60-90.html

OpenStudy (anonymous):

sides of a 30 60 90 triangle are always the same ratio

OpenStudy (anonymous):

ok this is what you do the ratios for a 30-60-90 right triangle are \(1:\sqrt{3}:2\) for short side: long side: hypotenuse you have long side = 7 so \[\frac{7}{\sqrt{3}}:7:\frac{14}{\sqrt{3}}\]